Write a recursive formula for the squence 8, 10, 12, 14, 16 then find the next term
Answer by math_helper(2461)

x[n+1] = x[n] + 2,     n=0,1,2,… ;  x[0] = 8

x[4] =  16
x[5] = x[4] + 2 = 16 + 2 = 18